OpenStudy (anonymous):

Use a graph of f or some other method to determine what, if any, value to assign to f(a) to make f continuous at x = a. f(x) = x/6x^2−x ; a=0

OpenStudy (dumbcow):

f(x) = x/x(6x-1) = 1/(6x-1) f(0) = -1